OpenStudy (anonymous):

show that the linearization of sin(x) at 0= the linearization of tan(x) at 0

OpenStudy (anonymous):

Linearization of sin(x) at x=0: L= f(0)-f'(0)*(x-0) L= 0 - cos(0*)x *the derivative of sin is cos L= -x Linearization of tan(x) at x=0 L= f(0)-f'(0)*x L=0-f'(sec^2(0))*x *the derivative of tan is sec^2 L=-x